# Nonequilibrium Probes of Quantum Geometry in Gapless Systems

### v1

@@ Abstract
Much of our understanding of gapless many-body quantum systems stems from their low-energy descriptions as conformal field theories. This is especially true in 1+1 dimensions, where such theories have an infinite-dimensional parameter space induced by their conformal symmetry. We reveal the associated quantum geometry by considering finite systems driven by time-dependent conformal transformations. For small deformations, perturbation theory predicts absorption rates and linear responses that are intrinsically related to components of the quantum geometric tensor. For arbitrarily large but adiabatic deformations, we show that periodic drives give rise to nontrivial return amplitudes involving the quantum metric, beyond the familiar leading order that only features a Berry phase. Our field-theoretic findings are universal, comprising general relations between measurable quantities and quantum geometry that only depend on the central charge of the conformal symmetry. This is supported by both analytical results for quantum dynamics under certain Floquet drives, and numerical simulations of gapless lattice models.
1 Introduction
Half a century ago, a seminal paper [1, 2] by Arnold put forward the idea that fluid flows trace geodesics in an infinite-dimensional manifold, revealing the key role played by infinite-dimensional (metric) geometry in physics [3, 4, 5]. The geometry in such cases is classical, related to the phase space of the system and to the kinetic energy in the classical Hamiltonian. In recent years though, it has come to be appreciated that quantum geometry—the geometry of bundles of quantum states over various parameter spaces [6, 7]—is essential to a number of phenomena in atomic and condensed matter physics; see e.g. [8, 9, 10, 11, 12, 13, 14, 15, 16] or the recent review [17] and references therein. This notably includes instances of quantum fluids and other many-body systems whose low-energy description involves an infinite-dimensional parameter space, blending Arnold’s ideas with those of modern quantum geometry.
A natural question, then, is whether infinite-dimensional quantum geometry can have observable consequences. That the answer must be affirmative is supported by the fact that any many-body quantum system is described, at low energy, by a quantum field theory, whose phase space is necessarily infinite-dimensional. Coherent states on that phase space define a bundle with an infinite-dimensional quantum geometry, as desired. For instance, Bloch coherent states for a single spin are labeled by two coordinates on a sphere, but their cousins for spin chains are labeled by infinitely many parameters (two at each lattice site) in the thermodynamic limit. While infinite-dimensional quantum geometry is ubiquitous in that sense, the issue is to find a setup that is observable, suitably universal, and sufficiently tractable even beyond small perturbations in parameter space. This work aims to provide just such a universal description for gapless systems, supported by both analytical computations and numerics.
@@ 1 Introduction
Figure 1: Return probability $|\langle\psi(0)|\psi(T)\rangle|^{2}$, starting from the initial ground state $|\psi(0)\rangle$ under the driven Hamiltonian (1.1) (extended to both right and left movers), plotted as a function of the (dimensionless) period $T/L$ for two different velocity profiles. (a) Velocity $v_{t}(x)=\cosh(2\lambda)-\sinh(2\lambda)\cos(2\pi k[x/L-t/T])$ for $\lambda=0.15$ and a single harmonic $k=2$. This profile only involves a finite-dimensional subalgebra of the Virasoro algebra, so the return probability can be evaluated exactly for any driving period (blue curve). (b) Velocity $v_{t}(x)=v\bigl[\cosh(2\lambda)-\sinh(2\lambda)\cos(2\pi k[x/L-t/T])+\frac{1}{5}\sin(2\pi k^{\prime}[x/L-t/T])\bigr]$ for $\lambda=0.15$, two harmonics $k=2$, $k^{\prime}=3$, and a normalization factor $v$. This profile involves the full infinite-dimensional Virasoro algebra, so no exact general expression is available for the return probability. In both cases, the return probability goes to 11 in the adiabatic limit $T/L\to\infty$. For finite $T/L$, this behavior is corrected by small oscillations whose amplitude is essentially the squared quantum distance between two nearby coherent states (dashed purple curves). When $T/L$ decreases, oscillations become more pronounced and are eventually no longer related to the quantum metric. The blue dots are the corresponding numerical simulations of the spin chain (1.2) illustrated in Fig. 2 with $\Delta=0$ and couplings $J_{j}(t)\propto v_{t}(x_{j})$ for $N=400$ sites $x_{j}$ (see Appendix E). Note the remarkable agreement between numerics and our analytical CFT results.
1.1 Driven conformal field theory
@@
Concretely, consider a chiral CFT with stress-energy tensor ${\mathcal{T}}(x)$, evolving under the time-dependent Hamiltonian [^1] Units are such that $\hbar=1$ and velocity is dimensionless, i.e. expressed in units of the ‘speed of light’.
$H(t)=\int_{0}^{L}\textrm{d}x\,v_{t}(x){\mathcal{T}}(x),$ (1.1)
@@
where the local velocity profile $v_{t}(x)>0$ varies smoothly in space and time. Space and time coordinates are respectively $x\in[0,L]$ and $t\in\mathbb{R}$, with $L$ the length (volume) of the system. We also impose periodicity in both space and time, namely ${\mathcal{T}}(x+L)={\mathcal{T}}(x)$, $v_{t}(x+L)=v_{t}(x)$, and $v_{t+T}(x)=v_{t}(x)$, where $T$ is the period of the drive. In standard CFT, the velocity profile would be homogeneous and constant: $v_{t}(x)=v=\textrm{const}$. By contrast, the Hamiltonian (1.1) is that of a chiral inhomogeneous CFT [21, 22, 23, 24, 25, 26] with a time-dependent velocity profile $v_{t}(x)$ generalizing earlier Floquet setups [27, 28, 29, 30, 31, 32, 33]. This extra, more general time-dependence will ultimately allow us to probe the quantum geometry of the system, as exemplified in Fig. 1.
@@
Concentrating on chiral CFTs with spatial periodicity is motivated not only by the simpler presentation, but also by its general applicability. Indeed, the setting is straightforwardly extended to the nonchiral case of decoupled right and left movers arising in many applications with periodic boundary conditions. An important example is provided by spin chains, whose low-energy excitations are described by the nonchiral version of (1.1). Think e.g. of an XXZ spin chain with $N$ sites and Hamiltonian
$H_{\textrm{XXZ}}(t)=-\sum_{j=1}^{N}J_{j}(t)\biggl[S^{x}_{j}S^{x}_{j+1}+S^{y}_{j}S^{y}_{j+1}-\Delta S^{z}_{j}S^{z}_{j+1}\biggr]\qquad(|\Delta|<1),$ (1.2)
@@
with $\mathfrak{su}(2)$ generators $S^{x,y,z}_{j}\equiv S^{x,y,z}_{j+N}$ in the spin-$1/2$ representation at site $j$, and time-dependent couplings $J_{j}(t)>0$ that vary on mesoscopic length scales222‘Mesoscopic’: on scales much larger than the lattice spacing $L/N$ but much smaller than the system size $L$. and satisfy $J_{j}(t)=J_{j}(t+T)$ and $J_{j}(t)=J_{j+N}(t)$. See Fig. 2 for an illustration. In addition, our setting can describe systems with open boundaries by folding/unfolding techniques adapted to inhomogeneous CFT [34, 35, 36].
Figure 2: Cartoon of a periodically driven gapless spin chain (1.2) of length $L\propto N$, with periodic boundary conditions. The inhomogeneous couplings $J_{j}(t)$ vary in space on mesoscopic length scales, and they vary smoothly in time with period $T$.
@@
Related questions on quantum geometry in CFT have recently been studied in holography [37, 38, 39, 40, 41] and quantum complexity [42, 43, 44]. Note, however, that many of the tools in this paper apply to a wide variety of quantum systems, well beyond CFT. Indeed, part of our work boils down to unitary representations of continuous groups, and the way these can be used to build time-dependent Hamiltonians. One instance is the family of Hamiltonians (1.1) for the Virasoro symmetry group of CFT, but many other physical systems fall into the same framework. Examples include the group $\mathrm{SU}(2)$ describing a spin in a rotating magnetic field [45], or the special linear group $\mathrm{SL}(2,\mathbb{R})$ describing both squeezed states in quantum optics [46] and Hall viscosity [47, 48]. In fact, a close cousin of the chiral theory (1.1) investigated here is the Hamiltonian of edge modes of quantum Hall droplets subjected to time-dependent area-preserving deformations [49, 50, 51].
Figure 3: The different regimes of interest in this work: drives in the perturbative regime are those for which the function $v_{t}(x)$ in Eq. (1.1) is close to being uniform, and drives in the adiabatic regime are those for which $v_{t}(x)$ oscillates slowly with $t$. When available, exact solutions cover the entire parameter space, but they are limited to specific families of deformations.
1.2 Summary of results
@@
In studying CFT Hamiltonians of the form (1.1), we obtain general results split into two categories: perturbative and adiabatic; see Fig. 3. These are universal in that they provide low-energy relations between measurable quantities and quantum geometry, which only depend on the central charge of the emergent effective description.
@@
  • On the perturbative end (Sec. 3), we consider velocities that only deviate slightly from a constant $v_{0}>0$, namely
@@
    where $\epsilon\ll 1$ and $w_{t}(x)$ is smooth but otherwise arbitrary. The resulting Hamiltonian (1.1) is close to that of a homogeneous CFT, so familiar formulas on time-dependent perturbation theory apply (provided $L$ is kept finite). In particular, picking $w_{t}(x)=\omega^{-1}\cos(\omega t)\partial_{x}X(x)$ with frequency $\omega>0$ and a time-independent function $X(x)$ yields an absorption rate $\Gamma(\omega)$ induced by transitions from the CFT ground state to excited states. We compute this rate and show, following [52], that its integral over frequencies is proportional to the quantum metric. In formulas,
    $\int_{0}^{\infty}\textrm{d}\omega\,\Gamma(\omega)\sim\frac{\epsilon^{2}}{2L^{2}}{\cal G}(\mathbf{X},\mathbf{X})$ (1.4)
    to leading order in $\epsilon$, where ${\cal G}(\cdot,\cdot)$ is the quantum metric evaluated at the identity in an infinite-dimensional space of inhomogeneous coherent states.
    Another natural quantity to investigate is the response of stress-energy expectation values to the perturbation (1.3). In that context, we pick an arbitrary (smooth) $L$-periodic function $Y(x)$, and ask how the time-dependent expectation value of ${\cal O}=\int\textrm{d}x\,Y(x){\mathcal{T}}(x)$ differs from its equilibrium value. We show that
    $\langle{\cal O}\rangle(t)-\langle{\cal O}\rangle\sim\epsilon\int_{0}^{t}\textrm{d}s\,\partial_{t}{\cal F}(\mathbf{Y}_{t-s},\mathbf{X}_{s})$ (1.5)
@@
  • On the adiabatic end (Sec. 4), we consider Hamiltonians (1.1) whose velocity profile $v_{t}(x)$ is well away from a constant, but varies slowly in time $t$ compared to the system size. We assume that the system is initially prepared in its ground state $|\psi(0)\rangle$, and investigate its overlap with the state $|\psi(T)\rangle$ at the end of an adiabatic cycle. To leading order in the adiabatic limit, the only difference between $|\psi(0)\rangle$ and $|\psi(T)\rangle$ is an overall phase—the sum of a dynamical phase and a Berry phase [53]. But, beyond leading order, there are new subleading corrections in the adiabatic parameter $\delta\sim L/T\ll 1$. Our interest lies in these corrections: we show that they probe the quantum metric of the space of conformal frames—the same metric as in the integrated absorption rate (1.4). In formulas, the return probability (Loschmidt echo) is
    $\bigl|\langle\psi(0)|\psi(T)\rangle\bigr|^{2}\sim 1-\delta^{2}{\cal G}(\mathbf{Y},\mathbf{Y}),$ (1.6)
@@
    where $\delta\mathbf{Y}$ is a small vector field measuring the deviation between the rays of $|\psi(T)\rangle$ and $|\psi(0)\rangle$. [^3] This vector field $\mathbf{Y}$ has nothing to do with the one in the linear response formula (1.5). We compute this deviation exactly for ‘rotating drives’ of the form $v_{t}(x)=v(x-\omega t)$, with a driving frequency $\omega=2\pi/T$ and a given profile $v(x)$ at $t=0$. The resulting return probability (1.6) gives rise to the oscillating pattern in Fig. 1.
@@
We stress that all these results are expansions, in different limiting regimes, of the same unified picture; see Fig. 3. However, access to the full range of drives is generally only possible via exact analytical computations for special classes of conformal transformations. These are studied in Sec. 5, and used there to benchmark our limiting results. An example is displayed in Fig. 1(a), showing excellent agreement between exact CFT return probabilities, their adiabatic expansions, and the corresponding numerical lattice computations.
1.3 Plan of the paper

### v2

@@ Abstract
Much of our understanding of gapless quantum matter stems from low-energy descriptions using conformal field theory. This is especially true in 1+1 dimensions, where such theories have an infinite-dimensional parameter space induced by their conformal symmetry. We reveal the underlying quantum geometry by considering finite many-body systems driven by time-dependent conformal transformations. For small deformations, perturbation theory predicts absorption rates and linear responses that probe the quantum geometric tensor. For arbitrarily large but adiabatic deformations, we show that periodic drives give rise to nontrivial return amplitudes involving the quantum metric, beyond the familiar leading order that only features a Berry phase. The former is less sensitive to decoherence than the latter, so it can provide robust experimental signatures of our predictions. Our field-theoretic findings are universal, comprising general relations between measurable quantities and quantum geometry that only depend on the emergent effective description. This is supported both by numerical simulations of gapless lattice models, and by exact results for quantum dynamics under certain Floquet drives, probing the full dynamical parameter space.
I Introduction
@@
Figure 1: Return probability $|\langle\psi(0)|\psi(T)\rangle|^{2}$ as a function of the period $T$, obtained by solving the Schrödinger equation for the nonchiral generalization (65) of the driven Hamiltonian (1). The initial state $|\psi(0)\rangle$ is the ground state of the initial Hamiltonian. The periodic drives in the two plots are given by (a) the single-harmonic velocity profile $v_{t}(x)=\cosh(2\lambda)-\sinh(2\lambda)\cos(2\pi k[x/L-t/T])$ with $\lambda=0.15$ and $k=2$, for which the return probability can be evaluated exactly for any driving period (blue curve, see Sec. V); (b) the velocity $v_{t}(x)=v\bigl[\cosh(2\lambda)-\sinh(2\lambda)\cos(2\pi k[x/L-t/T])+\frac{1}{5}\sin(2\pi k^{\prime}[x/L-t/T])\bigr]$ with $\lambda=0.15$, two harmonics $k=2$, $k^{\prime}=3$, and a normalization $v$. In both cases, the return probability goes to 11 in the adiabatic limit $T/L\to\infty$. For finite $T/L$, this behavior is corrected by oscillations whose amplitude is essentially the squared quantum distance between two nearby coherent states, which we compute explicitly (dashed red curve, see Sec. IV). The blue dots are the corresponding numerical simulations of the spin chain (2) illustrated in Fig. 2, with $\Delta=0$ and couplings $J_{j}(t)$ of the same form as $v_{t}(x_{j})$ for $N=L=402$ sites $x_{j}$ (see Appendix E). Note the remarkable agreement between numerics and our analytical CFT results.
Half a century ago, a seminal paper by Arnold [1] put forward the idea that fluid flows trace geodesics in an infinite-dimensional manifold, revealing the key role played by infinite-dimensional (metric) geometry in physics [2, 3, 4]. The geometry in such cases is classical, related to the phase space and the kinetic energy of the system. In recent years though, it has come to be appreciated that quantum geometry—the geometry of bundles of quantum states over parameter spaces [5, 6]—is essential for condensed matter physics [7, 8]. Applications range e.g. from band theory [9, 10, 11] to strongly correlated electrons [12, 13, 14, 15, 16, 17, 18, 19], through quantum phase transitions [20], time-dependent variational principles [21, 22, 23] and Floquet engineering [24, 25].
A natural question is whether Arnold’s geometric insight on fluids has a quantum counterpart, and whether it has observable implications in the context of quantum geometry. That the answer should be positive is supported by the fact that any many-body quantum system is described, at low energies, by a quantum field theory, whose phase space is necessarily infinite-dimensional. Coherent states on that phase space define a bundle with an infinite-dimensional quantum geometry, as desired. For instance, Bloch coherent states for a single spin are labeled by two coordinates on a sphere, but their cousins for spin chains are labeled by infinitely many parameters (two at each lattice site) in the thermodynamic limit. While infinite-dimensional quantum geometry is ubiquitous in that sense, the nontrivial issue is to find a setup that is observable, suitably universal, and sufficiently tractable even beyond small perturbations in parameter space.
This work aims to provide just such a universal description for gapless quantum matter, supported by both analytical computations and numerics. Specifically, we consider nonequilibrium setups that probe the quantum metric and the Berry curvature: the former measures the distance between pure states [5], while the latter quantifies phase variations in parameter space [26]. These quantities are what is meant in practice by quantum geometry, and knowing them provides a handle on the dauntingly vast space of quantum states. In particular, the metric is expected to be less sensitive to decoherence than Berry phases, thereby providing robust experimental signatures to put our predictions to the test.
Figure 2: Cartoon of a periodically driven gapless spin chain (2) of length $L\propto N$ with periodic boundary conditions. The inhomogeneous couplings $J_{j}(t)$ vary on mesoscopic scales in space, and continuously in time with period $T$.
I.1 Driven conformal field theory
@@
Concretely, consider a right-moving chiral CFT with stress-energy tensor ${\mathcal{T}}(x)$, evolving under the time-dependent Hamiltonian [^1] Units are such that $\hbar=1$ and velocity is dimensionless, i.e. expressed in units of the ‘speed of light’.
$H(t)=\int_{0}^{L}\textrm{d}x\,v_{t}(x){\mathcal{T}}(x).$ (1)
@@
Here the velocity profile $v_{t}(x)>0$ depends continuously on position $x\in[0,L]$ and time $t\in\mathbb{R}$, with $L$ the length (volume) of the system. We impose periodicity in space, namely ${\mathcal{T}}(x+L)={\mathcal{T}}(x)$ and $v_{t}(x+L)=v_{t}(x)$, and often also in time so that $v_{t+T}(x)=v_{t}(x)$, with $T$ then the period of the drive. In standard CFT, the velocity would be homogeneous and constant: $v_{t}(x)=v=\textrm{const}$. By contrast, the Hamiltonian (1) is deformed to that of a chiral inhomogeneous CFT [30, 31, 32, 33, 34, 35] with a time-dependent velocity profile $v_{t}(x)$ generalizing earlier Floquet setups [36, 37, 38, 39, 40, 41, 42]. This extra, more general time-dependence will ultimately allow us to probe the quantum geometry of the system, as exemplified in Fig. 1.
@@
Concentrating on chiral CFTs with spatial periodicity is motivated not only by the simpler presentation, but also by its general applicability. Indeed, the setting is straightforwardly extended to the nonchiral case of decoupled right and left movers arising in many applications with periodic boundary conditions. An important example is provided by driven inhomogeneous spin chains, whose low-energy excitations are described by the nonchiral version of (1). Think e.g. of an $XXZ$ spin chain with $N$ sites and Hamiltonian
$H_{\textrm{XXZ}}(t)=-\sum_{j=1}^{N}J_{j}(t)\Bigl(S^{x}_{j}S^{x}_{j+1}+S^{y}_{j}S^{y}_{j+1}-\Delta S^{z}_{j}S^{z}_{j+1}\Bigr)$ (2)
@@
with $|\Delta|<1$ and spin-$1/2$ $\mathfrak{su}(2)$ generators $S^{x,y,z}_{j}\equiv S^{x,y,z}_{j+N}$ at site $j$. The $J_{j}(t)>0$ are continuously driven couplings that vary on mesoscopic scales in space222‘Mesoscopic’: on length scales much larger than the lattice spacing $L/N$ but much smaller than the system size $L$. and satisfy $J_{j+N}(t)=J_{j}(t)=J_{j}(t+T)$. See Fig. 2 for an illustration. In addition, our setup can describe systems with open boundaries by folding/unfolding techniques adapted to inhomogeneous CFT [43, 44, 45].
Related questions on quantum geometry in CFT have been studied in holography [46, 47, 48, 49, 50, 51, 52] and quantum complexity [53, 54, 55], following earlier works on Virasoro coadjoint orbits [56, 57, 58, 59]. [^3] There also exists a metric on the space of marginal deformations of 1+1D CFTs [60], but this is unrelated to our framework since (1) is not a marginal deformation of a homogeneous Hamiltonian. Models such as (1) have also been used to investigate many-body quantum quenches from initial states with inhomogeneous temperature profiles [61, 33, 35]. In such cases, the interpretation of $v_{t}(x)$ as a local (inverse) temperature links thermal transport to metric deformations [62, 33, 35, 63] and provides a smooth version of gluing two systems at different temperatures [64, 65, 66].
@@
Note, however, that many of the tools in this paper apply to a wide variety of quantum systems, well beyond CFT. Indeed, part of our work boils down to unitary representations of continuous groups, and the way these can be used to build time-dependent Hamiltonians. One instance is the family of Hamiltonians (1) for the Virasoro symmetry group of CFT, but many other physical systems fall into the same framework. Examples include the group $\mathrm{SU}(2)$ describing a spin in a rotating magnetic field [26], or the special linear group $\mathrm{SL}(2,\mathbb{R})$ describing both squeezed states in quantum optics [67] and Hall viscosity [68, 69]. In fact, a close cousin of the chiral theory (1) investigated here is the Hamiltonian of edge modes of quantum Hall droplets subjected to time-dependent area-preserving deformations [70, 71, 72].
I.2 Summary of results
@@
In studying CFT Hamiltonians of the form (1), we obtain general results split into two categories: perturbative and adiabatic; see Fig. 3. These are universal in that they relate quantum geometry to measurable quantities at low energies, in a way that only depends on the central charge of the emergent effective theory.
Perturbative.
@@
On the perturbative end (Sec. III), we consider velocities that only deviate slightly from a constant $v_{0}>0$, namely
@@
where $\epsilon\ll 1$ and $w_{t}(x)=w_{t}(x+L)$ is smooth but otherwise arbitrary. The resulting Hamiltonian (1) is close to that of a homogeneous CFT, so familiar formulas on time-dependent perturbation theory apply (provided $L$ is kept finite). In particular, picking $w_{t}(x)=(\omega L)^{-1}\cos(\omega t)\partial_{x}X(x)$ with frequency $\omega>0$ and a time-independent function $X(x)$ yields an absorption rate $\Gamma(\omega)$ induced by transitions from the CFT ground state to excited states. [^4] We denote vector fields on the circle as $\mathbf{X}=X(x)\partial_{x}$, and similarly for $\mathbf{Y}$ and $\mathbf{W}$. We compute this rate and show, following [73], that its integral over frequencies is proportional to the quantum metric. In formulas,
$\int_{0}^{\infty}\textrm{d}\omega\,\Gamma(\omega)\sim\frac{\epsilon^{2}}{2L^{2}}\,{\cal G}(\mathbf{X},\mathbf{X})$ (4)
to leading order in $\epsilon$, where ${\cal G}(\cdot,\cdot)$ is the quantum metric evaluated at the identity in an infinite-dimensional space of inhomogeneous coherent states. This link mirrors the fluctuation-dissipation theorem, and is similar in that sense to the defining relation of quantized dichroism for quantum Hall systems [74, 75, 76, 77]. Remarkably, it is possible to give explicit expressions for the quantum metric despite the infinite-dimensional parameter space [see Eqs. (29a)–(30a) and (32)].
Figure 3: The different regimes of interest in this work: drives in the perturbative regime are those for which the function $v_{t}(x)$ in Eq. (1) is close to being uniform, and drives in the adiabatic regime are those for which $v_{t}(x)$ oscillates slowly with $t$. When available, exact solutions cover the entire parameter space, but they are limited to specific families of deformations.
Another natural quantity to investigate is the response of stress-energy expectation values to the perturbation (3), using $w_{t}(x)=\partial_{x}X_{t}(x)$ for an arbitrary time-dependent profile $X_{t}(x)$. In that context, we pick an arbitrary (smooth) $L$-periodic function $W(x)$, and ask how the time-dependent expectation value of ${\cal O}=\int\textrm{d}x\,W(x){\mathcal{T}}(x)$ differs from its equilibrium value. We show that
$\langle{\cal O}\rangle(t)-\langle{\cal O}\rangle\sim\epsilon\int_{0}^{t}\textrm{d}s\,{\cal F}(\partial_{s}\mathbf{W}_{t-s},\mathbf{X}_{s})$ (5)
@@
On the adiabatic end (Sec. IV), we consider Hamiltonians (1) whose velocity profile $v_{t}(x)$ is well away from a constant, but varies slowly in time compared to the system size. We assume that the system is initially prepared in its ground state $|\psi(0)\rangle$, and investigate its overlap with the state $|\psi(T)\rangle$ at the end of an adiabatic cycle. To leading order in the adiabatic limit, the only difference between $|\psi(0)\rangle$ and $|\psi(T)\rangle$ is an overall phase—the sum of a dynamical phase and a Berry phase [46]. But, beyond leading order, there are new subleading corrections in the adiabatic parameter $\delta\sim L/T\ll 1$. Our interest lies in these corrections: they probe the quantum metric of the space of conformal frames—the same metric as in the integrated absorption rate (4). In formulas, the return probability (Loschmidt echo) is
$\bigl|\langle\psi(0)|\psi(T)\rangle\bigr|^{2}\sim 1-\delta^{2}{\cal G}(\mathbf{Y}_{T},\mathbf{Y}_{T}),$ (6)
@@
where $\delta\mathbf{Y}_{T}$ is a small vector field measuring the deviation between the rays of $|\psi(T)\rangle$ and $|\psi(0)\rangle$. We compute this deviation exactly for ‘rotating drives’ of the form $v_{t}(x)=v(x-\omega t)$, with a driving frequency $\omega=2\pi/T$ and a given profile $v(x)$ at $t=0$. The resulting return probability (6) gives rise to the oscillating pattern in Fig. 1.
It should not come as a surprise that adiabatic overlaps such as (6) probe the quantum metric. After all, the latter is the pullback of the Fubini-Study metric that measures, by definition, the overlap of rays in projective Hilbert space. What is special in our case is that we can compute such overlaps dynamically, at any time $t$, by providing an approximate solution $|\psi(t)\rangle$ of the Schrödinger equation for a time-dependent CFT Hamiltonian (1). More precisely, we compute the time-dependent deviation $\mathbf{Y}_{t}$ between $|\psi(t)\rangle$ and its leading-order adiabatic approximation, whereupon Eq. (6) follows at $t=T$. A key aspect of this derivation is to identify the ‘correct’ adiabatic parameter $\delta$, thanks to which the adiabatic expansion of $|\psi(t)\rangle$ can in principle be pushed to arbitrarily high order.
Beyond expansions.
@@
We stress that all the results above are expansions, in different limiting regimes, of the same unified picture; see Fig. 3. However, access to the full range of drives is generally only possible via exact analytical computations for special classes of conformal transformations. These are studied in Sec. V, and used there to benchmark our limiting results. An example is displayed in Fig. 1(a), showing excellent agreement between exact CFT return probabilities, their adiabatic expansions, and the corresponding numerical lattice computations.
I.3 Plan of the paper